OpenStudy (otakulyfe):

What is the equation of the line that passes through (1, 3) and (-2, -3)?

OpenStudy (otakulyfe):

@mathmate

OpenStudy (mathmate):

There is a cumbersome formula that allows you to put in the four numbers somewhere in the formula and get the equation instantly. But most people do it in two simple steps: Assume P1(1,3),P2(-2,-3) 1. find the slope of the line using \(slope, m=\Large \frac{y_2-y_1}{x_2-x_1}\) 2. use the point slope form of the equation of a line: \(\large (y-y_1)=m(x-x_1)\) Simplify or beautify as necessary.

OpenStudy (mathmate):

Example: line between P1(1,2), P2(5,4) m=(4-2)/(5-1)=2/4=(1/2) equation of line (y-y1)=m(x-x1) y-2=(1/2)(x-1) y=(1/2)x -1/2+2 y=(1/2)x+ 3/2 ...........equation of line passing through P1(1,2), P2(5,4)

OpenStudy (otakulyfe):

thanks

OpenStudy (mathmate):

You're welcome! :)